Publications

Forrest M. Hoffman and William W. Hargrove. Cluster computing: Linux taken to the extreme. Linux

Magazine, 1(1):56–59, 1999. Forrest M. Hoffman. Concepts in Beowulfery. Linux Magazine, 4(1):40–41, January 2002a. Forrest M. Hoffman. Configuring a Beowulf Cluster. Linux Magazine, 4(2):42–45, February 2002b. Forrest M. Hoffman. An Introduction to Parallel Programming. Linux Magazine, 4(3):38–41, March 2002c. Forrest M. Hoffman. Message Passing with MPI and PVM. Linux Magazine, 4(4):38–41, 63, April 2002d. Forrest M. Hoffman. Advanced Message Passsing. Linux Magazine, 4(5):42–43, 62, May 2002e. Forrest M. Hoffman. Using Mosix. Linux Magazine, 4(6):46–49, June 2002f. Forrest M. Hoffman. Scalable I/O on Clusters, Part 1. Linux Magazine, 4(7):42–43, 59, July 2002g. Forrest M. Hoffman. Scalable I/O on Clusters, Part 2. Linux Magazine, 4(8):48–49, 59, August 2002h. Forrest M. Hoffman. High Performance Interconnects. Linux Magazine, 4(9):46–47, September 2002i. Forrest M. Hoffman. Job Scheduling and Batch Systems. Linux Magazine, 4(10):40–43, October 2002j. Forrest M. Hoffman. Job Scheduling with Maui. Linux Magazine, 4(11):44–47, November 2002k. Forrest M. Hoffman. Running Commands on Many Nodes. Linux Magazine, 4(12):42–45, December 2002l. Forrest M. Hoffman. Kickstarting Cluster Nodes, Part 1. Linux Magazine, 5(1):46–49, 63, January 2003a. Forrest M. Hoffman. Kickstarting Compute Nodes, Part 2. Linux Magazine, 5(2):44, 46–49, 63, February

2003b. Forrest M. Hoffman. Master/Slave Message Passing. Linux Magazine, 5(3):50, 52, 54–55, 61, March 2003c. Forrest M. Hoffman. Using Derived Data Types with MPI. Linux Magazine, 5(4):42, 44–47, April 2003d. Forrest M. Hoffman. MPI Communicators and Groups. Linux Magazine, 5(5):42–43, 46–47, 67, May 2003e. Forrest M. Hoffman. Coupling Paralllel Models. Linux Magazine, 5(6):36, 38, 40–44, 46–47, 71, June 2003f. Forrest M. Hoffman. Communicators & Groups, Part 2. Linux Magazine, 5(7):46, 48–51, July 2003g. Forrest M. Hoffman. Cluster Monitoring with Ganglia. Linux Magazine, 5(8):42, 44, 46–47, August 2003h. Forrest M. Hoffman. It’s All About Speed. Linux Magazine, 5(9):44–47, September 2003i. Forrest M. Hoffman. Management with Condor, Part 1. Linux Magazine, 5(10):42–44, 62, October 2003j. Forrest M. Hoffman. Management with Condor, Part 2. Linux Magazine, 5(11):42–44, November 2003k. Forrest M. Hoffman. Management with Condor, Part 3. Linux Magazine, 5(12):34–37, 59, December 2003l. Forrest M. Hoffman. Multi-Processing with OpenMP. Linux Magazine, 6(1):44–48, January 2004a. Forrest M. Hoffman. OpenMP Multi-Processing, Part 2. Linux Magazine, 6(2):42–46, February 2004b. Forrest M. Hoffman. Using OpenMP, Part 3. Linux Magazine, 6(3):44–48, March 2004c. Forrest M. Hoffman. Writing Hybrid MPI/OpenMP Code. Linux Magazine, 6(4):42–45, 63, April 2004d. Forrest M. Hoffman. 64-bit Computing with SGI’s Altix. Linux Magazine, 6(5):40–43, May 2004e.

Forrest M. Hoffman. Parallel netCDF. Linux Magazine, 6(7):55–58, July 2004f. Forrest M. Hoffman. The Future of Message Passing. Linux Magazine, 6(8):54–56, 63, August 2004g. Forrest M. Hoffman. Using MPI-2. Linux Magazine, 6(9):54–57, September 2004h. Forrest M. Hoffman. One-Sided Communications with MPI-2. Linux Magazine, 6(11):54–58, 60, 62, Novem-

ber 2004i. Forrest M. Hoffman. A Very Super Computing Show. Linux Magazine, 7(1):54–56, January 2005a. Forrest M. Hoffman. Using BProc. Linux Magazine, 7(2):52, 54–55, February 2005b. Forrest M. Hoffman. Building a (Very) Low Cost Cluster. Linux Magazine, 7(3):54–55, 63, March 2005c. Forrest M. Hoffman. Clustermatic. Linux Magazine, 7(4):39–40, 42, 44–45, April 2005d. Forrest M. Hoffman. The Rocks Cluster Distribution. Linux Magazine, 7(5):45–57, May 2005e. Forrest M. Hoffman. Computing with OSCAR. Linux Magazine, 7(7):48–50, 61, July 2005f. Forrest M. Hoffman. Scyld Beowulf. Linux Magazine, 7(8):52, 54, 56, 70, August 2005g. Forrest M. Hoffman. Designing a Cluster Solution. Linux Magazine, 7(9):56–57, 66, September 2005h. Forrest M. Hoffman. HPC Logging with syslog-ng, Part One. Linux Magazine, 7(11), November 2005i. Forrest M. Hoffman. HPC Logging with syslog-ng, Part Two. Linux Magazine, 8(1):50–52, 63, January

2006a. Forrest M. Hoffman. A Conference for Extremophiles. Linux Magazine, 8(2):54–55, 63, February 2006b. Forrest M. Hoffman. Unified Parallel C 101. Linux Magazine, 8(3):48–51, March 2006c. Forrest M. Hoffman. Unified Parallel C, Part Two. Linux Magazine, 8(5):48–51, May 2006d. Forrest M. Hoffman. Unified Parallel C, Part Three. Linux Magazine, 8(6):48–50, 57, June 2006e. Forrest M. Hoffman. The Global Arrays Toolkit. Linux Magazine, 8(7):42–44, July 2006f. Forrest M. Hoffman. Global Arrays Toolkit, Part Two. Linux Magazine, 8(8):48–51, 59, August 2006g. Forrest M. Hoffman. Global Arrays Toolkit, Part Three. Linux Magazine, 8(9):40–43, 53, September 2006h. Forrest M. Hoffman. The Future of HPC. Linux Magazine, 8(10):46–48, October 2006i.